OpenStudy (3psilon):

Lim x->0 sinx/x^2+4x why is it 1/4?!

hero (hero):

Hint: sin(x)/x --> 1

OpenStudy (3psilon):

I knew it had to do with the squeeze theorem that x^2 is throwing me off

OpenStudy (anonymous):

We can use a simple identity to solve such: \[\lim_{x\to0}\left(\frac{\sin x}{x^2+4x}\right)=\lim_{x\to0}\left(\frac{\sin x}{x(x+4)}\right)=\lim_{x\to0}\left(\frac{1}{x+4}\right)=\frac{1}{4}\]

OpenStudy (3psilon):

AHhhh I see it!

OpenStudy (anonymous):

(Sorry, I was about to do l'Hospital's and then I saw the factorization at the bottom, @Hero !)

OpenStudy (3psilon):

Thanks @LolWolf

hero (hero):

I suppose my hint was useless